OpenStudy (anonymous):

Using the given points, determine the slope. (0, 32) and (100, 212) slope = -9/5 slope = -5/9 slope = 5/9 slope = 9/5

OpenStudy (tkhunny):

Have you considered the Slope Formula? Change in y is what?? 212-32 Change in x is what?? 100-0

OpenStudy (anonymous):

I did that and didn't get an answer

OpenStudy (tkhunny):

That makes no sense. How can you get no answer. 212-32 = 180 100-0 = 100 180/100 = 18/10 = 9/5 = Slope. So, what was it that you did?